How many grams of water vapor (H2O) are in a 10.2 liter sample at 0.98 atmospheres and 26°C? Show all work used to solve this problem.
1
Expert's answer
2013-05-15T11:01:18-0400
For solving that task we shall use the equation:
pV = nRT,
where p-pressure, V-volume, n - amount of moles, n - gas constant (n = 0.080206)
n = pV/RT
The amount of moles can be found by the equation:
n = m/M
where m - mass, M - molar mass
So the mass could be calculated as:
m = n*M
M(H2O) = 18 grams/mole
n(H2O) = 0.98*10.2/0.08206*(26 + 273) = 0.43 moles
m(H2O) = 0.43*18 = 7.3 grams